# Computationally-Efficient Linear Periodically Time-Variant Digital PLL Modeling Using Conversion Matrices and Uncorrelated Upsampling

### v2

@@ IV-E Algorithm Complexity
Figure 15: Computation time for one frequency-domain sample with different problem dimensions. The green dots are actual computation time and the fitted curves are shown in black.
V Conclusion
@@
This Appendix provides the intuition that helps finding the sequence $w_{U,L}[n]$. In Fig. 5, $x[m]$ is a WSS low sampling rate sequence that needs to be upsampled. The DTFT of $x[m]$, $X(\Omega)$, is shown in Fig. 16(a). The height inside the shaded area at a specific $\Omega$ contains information on correlation; having different shaded height means uncorrelated spectrum. For example, $X(a)$ is uncorrelated with $X(b)$, i.e., $E\{X(a)X(b)\}=0$, because they have different shaded height. On the other hand, $X(a)$ is correlated with $X(c)$, i.e., $E\{X(a)X(c)\}\neq 0$, for having the same shaded height. When $x[m]$ is a real WSS random process, $X(\Omega)$ is conjugate symmetric w.r.t. $\Omega=k\pi$ and aperiodic for $\Omega\in[0,2\pi)$.
After upsampling by $L$ times by inserting $L-1$ zeros between adjacent samples, $y_{corr}[n]$ is no longer a WSS random process, but a cyclostationary random process. The autocorrelation sequence of $y_{corr}[n]$, $R_{yy,corr}$, is $L$-periodic, making the period of $Y_{corr}(\Omega)$ equal to $2\pi/L$. Fig. 16 shows $Y_{corr}(\Omega)$ when $L=6$.
When $y_{corr}[n]$ passes through a system with fundamental frequency $2\pi/N$, the output spectrum of the system $Y_{sys}(\Omega)$ is a summation of frequency-shifted $Y_{corr}(\Omega)$:
$Y_{sys}(\Omega)=\sum_{k=0}^{N-1}H_{0,k}(\Omega)Y_{corr}\left(\Omega-\frac{2\pi k}{N}\right).$ (26)
Fig. 16(c) shows $Y_{corr}(\Omega-\pi/4)$ for $N=8$ and $k=1$. By inspection, $Y_{corr}(\Omega-\pi/4)$ is uncorrelated with $Y_{corr}(\Omega)$. In general, there are 2 types of frequency shifts:
@@
    Correlated shift: the shifted spectrum is correlated with the original spectrum. For $L=6$ and $N=8$, $k_{cs}=0,4$.
@@
    Uncorrelated shift: the shifted spectrum is uncorrelated with the original spectrum. For $L=6$ and $N=8$, $k_{us}=1,2,3,5,6,7$.
Since (5) will be used to calculate the output PSD, $Y_{corr}(\Omega)$ needs to be pre-decorrelated before feeding into the system. By inspection, the sequence of correlated shift is
$k_{cs}[n]=\begin{cases}1,&\text{if mod}(n,N/\mathrm{gcd}(L,N))=0\\ {0},&{\text{else}}\end{cases}.$ (27)

### v3

@@ IV-F Cross Verification
Figure 23: (a) Computation time for one frequency-domain sample with different problem dimensions. The green dots are the actual computation time, and the fitted curves are shown in black. (b) Comparison of computation time for $2^{14}$ frequency-domain samples between the proposed work and the simulation. The number of periodogram averages is set to 500 for the simulation.
IV-G Algorithm Complexity
Fig. 23(a) shows the computation time required for a single frequency-domain sample using the proposed conversion matrix-based algorithm across varying problem dimensions $N$, specifically when $M=N-1$ and $P=2$. This algorithm was evaluated on a laptop equipped with an AMD Ryzen 9 5900HS CPU and 16.0 GB of RAM. For values where $N<150$, computation time exhibits a relationship proportional to $N^{2}$. However, for greater dimensions ($N\geq 200$), the complexity appears to be proportional to $N^{2.5}$. The proposed uncorrelated upsampling approach can reduce computation time by a factor of at least $(N-1)^{2}$. When $N<65$, the computation time for one frequency-domain sample remains under 1 ms, and consumes less time than the simulation for $N<70$ as shown in Fig. 23(b). The efficiency of the algorithm can be further leveraged if only a few, instead of $2^{14}$, spot PN needs to be optimized, allowing the designer to quickly explore the design space to fine-tune PN at selected frequencies and study noise folding with great accuracy.
V Limitations
The proposed work has the following limitations:
  1. 1.
    Linear circuits. In the case of a bang-bang PLL, this work can serve as a analysis tool to study noise folding, with the linear BBPD/OT-TDC gain extracted from either a behavioral transient simulation or a time-domain jitter analysis [15, 6, 44].
  2. 2.
    Integer-$N$ PLLs. Fractional-$N$ operation makes $f_{\text{DCO}}$ a non-integer multiple of $f_{\text{REF}}$, which cannot be easily captured by the proposed model.
VI Conclusion
@@
This Appendix provides the intuition that helps finding the sequence $w_{\text{U},L}[n]$. In Fig. 6, $x[m]$ is a WSS low sampling rate sequence that needs to be upsampled. The DTFT of $x[m]$, $X(\Omega)$, is shown in Fig. 24(a). The height inside the shaded area at a specific $\Omega$ contains information on correlation; having different shaded height means uncorrelated spectrum. For example, $X(a)$ is uncorrelated with $X(b)$, i.e., $E\{X(a)X(b)\}=0$, because they have different shaded height. On the other hand, $X(a)$ is correlated with $X(c)$, i.e., $E\{X(a)X(c)\}\neq 0$, for having the same shaded height. These shades are only for illustrating the correlation, not the actual value of these spectra. When $x[m]$ is a real WSS random process, $X(\Omega)$ is conjugate symmetric w.r.t. $\Omega=k\pi$ and aperiodic for $\Omega\in[0,2\pi)$.
After upsampling by $L$ times by inserting $L-1$ zeros between adjacent samples, $y_{\text{corr}}[n]$ is no longer a WSS random process, but a cyclostationary random process. The autocorrelation sequence of $y_{\text{corr}}[n]$, $R_{yy,\text{corr}}$, is $L$-periodic, making the period of $Y_{\text{corr}}(\Omega)$ equal to $2\pi/L$. Fig. 24 shows $Y_{\text{corr}}(\Omega)$ when $L=6$.
When $y_{\text{corr}}[n]$ passes through a system with fundamental frequency $2\pi/N$, the output spectrum of the system $Y_{\text{sys}}(\Omega)$ is a summation of frequency-shifted $Y_{\text{corr}}(\Omega)$:
$Y_{\text{sys}}(\Omega)=\sum_{k=0}^{N-1}H_{0,k}(\Omega)Y_{\text{corr}}\left(\Omega-\frac{2\pi k}{N}\right).$ (28)
Figure 25: Example of the DSM conversion matrix $\mathbf{H}_{\text{DSM,tot}}$ at 10 kHz for $N=6$ and $M=2$ and its decorrelation conversion matrix $\mathbf{W}_{\text{U},M}$.
Fig. 24(c) shows $Y_{\text{corr}}(\Omega-\pi/4)$ for $N=8$ and $k=1$. By inspection, $Y_{\text{corr}}(\Omega-\pi/4)$ is uncorrelated with $Y_{\text{corr}}(\Omega)$. In general, there are 2 types of frequency shifts:
@@
    Correlated shift: the shifted spectrum is correlated with the original spectrum. For $L=6$ and $N=8$, $k_{\text{cs}}=0,4$.
@@
    Uncorrelated shift: the shifted spectrum is uncorrelated with the original spectrum. For $L=6$ and $N=8$, $k_{\text{us}}=1,2,3,5,6,7$.
Since (5) will be used to calculate the output PSD, $Y_{\text{corr}}(\Omega)$ needs to be pre-decorrelated before feeding into the system. By inspection, the sequence of correlated shift is
$k_{\text{cs}}[n]=\begin{cases}1,&\text{if mod}(n,N/\mathrm{gcd}(L,N))=0\\ {0},&{\text{else}}\end{cases}.$ (29)
@@
Fig. 25 shows the conversion matrix of the decorrelation sequence $\mathbf{W}_{\text{U},M}$ for $N=6$ and $M=2$. All non-zero elements are separated by 3 circularly in columns and rows, which equals to the period of the correlated shift sequence $k_{\text{cs}}[n]$. Multiplying $\mathbf{W}_{\text{U},M}$ on the right side of the DSM conversion matrix before decorrelation $\mathbf{H}_{\text{corr}}$ provides a correction factor on its bottom row, which compensates the underestimated conversion gain caused by ignoring cross-correlated spectrum and using time-averaged PSD of a cyclostationary input.
Acknowledgments
This work is supported by Semiconductor Research Corporation Task No. 3160.040 through UT Dallas’ Texas Analog Center of Excellence (TxACE).